# Trajectory learning for ensemble forecasts via the continuous ranked probability score: a Lorenz ’96 case study

### v1

@@ 4 Model performance
The weather forecasts are carried out via ensemble predictions for 2 MTU, where each ensemble comprising 50 members. The slow variables sampled from 100 high-fidelity solution snapshots are used as initial conditions. These snapshots, each separated by 10 MTU, lie outside of the training/validation dataset. We consider perfectly known initial conditions in Section 4.1 and perturbed initial conditions in Section 4.3. The prediction ensembles are assessed via the mean squared error (MSE), the spread-error relationship, and the CRPS. The predicted climate is assessed via a single long-time ensemble prediction, computed over 3000 MTU. The resulting distributions of the slow variables are compared via the Kolmogorov–Smirnov distance and the Hellinger distance. All error assessment metrics are further detailed in Appendix B. A summary of the models tested in this section is given in Table 3.
@@ 4.1 Short-time predictions
A further increase in average accuracy is observed for the CRPS-based models. Increasing the trajectory length generally leads to parametrizations with improved accuracy, up to a certain limit. For example, the additive noise model trained at $N_{t}=8$ is more accurate than the model trained at $N_{t}=2$, and is approximately as accurate as that trained at $N_{t}=32$. The multiplicative models exhibit an even smaller MSE regardless of the trajectory length used in training, suggesting that the additional degrees of freedom allow for increased prediction accuracy.
@@
For the trained models, it could be desirable to diminish or increase the weight of the self-spread term in the CRPS loss to favor statistical consistency. This can possibly be achieved by means of a hyperparameter, or by introducing a term penalizing the discrepancy between the ensemble spread and MSE. Such a modification can diminish the ensemble sharpness during training without sacrificing accuracy.
@@ 4.2 Long-time predictions
The climate prediction using the stochastic parametrizations is shown in Fig. 7, defined via the empirical PDFs of the slow variables over a long simulation time (3000 MTU). Here, the ‘crps_mult8’ is not included in the results, as it exhibited numerical instabilities during extended simulation. Generally, all stochastic parametrizations produce qualitatively similar symmetric PDFs with an excessively large standard deviation. An improvement of the CRPS-based models is found compared to the derivative-fitting-based models, which becomes clearer when increasing the trajectory length. This is summarized in Table 4, indicating that the additive noise models learned using longer trajectory lengths outperform the other tested models. However, the results are not as accurate as those presented in literature (see, e.g., [16], [10]), suggesting that combining CRPS-based learning with other state-dependent model forms might be beneficial for long-time prediction accuracy.
@@ 5 Conclusion
In this paper, we investigated the calibration of stochastic parametrizations for ensemble forecasting using trajectory learning based on the continuous ranked probability score (CRPS). Using the two-scale Lorenz ’96 system as a testbed, we demonstrated that this approach yields models that provide both accurate and sharp short-term forecasts. Both additive and multiplicative noise models, leveraging global spatial patterns from POD modes, were found to outperform models based on derivative fitting. The improved accuracy at short lead times suggests that CRPS-based trajectory fitting is promising approach when combined with data assimilation, where accurate short-term forecasts are essential. Ultimately, this work establishes CRPS-based trajectory learning as a promising methodology for developing data-driven stochastic parametrizations for complex dynamical systems, representing a first step towards its application in geophysical flows.
@@ Appendix A Model and training details
The parameters for the derivative-fitting-based models with correlated noise, introduced in Section 3.2, are summarized in Table 5. These models use increments generated by an Ornstein–Uhlenbeck process, discretized as a first-order autoregressive (AR(1)) model, for the stochastic model component. The specific autocorrelation coefficients per POD mode are provided. Fig. 9 provides a visual comparison between the measured autocorrelation from the data and the autocorrelation function of the AR(1)-processes with fitted parameters. As discussed in Section 4.1, this figure highlights a potential overestimation of the temporal correlation.
@@ Appendix B Forecast assessment metrics
In the following definitions, we denote the truth by $y$ and denote by $x$ an ensemble forecast of size $M$, consisting of ensemble members $x_{i},\leavevmode\nobreak\ i=1,\ldots,M$. All $x_{j}$ and $y$ are regarded as independent identically distributed random variables with mean $\mu$ and variance $\sigma^{2}$.

### v2

@@ 4 Model performance
The weather forecasts are carried out via ensemble predictions for 2 MTU, where each ensemble comprises 50 members. The slow variables sampled from 100 high-fidelity solution snapshots are used as initial conditions. These snapshots, each separated by 10 MTU, lie outside of the training/validation dataset. We consider perfectly known initial conditions in Section 4.1 and perturbed initial conditions in Section 4.3. The prediction ensembles are assessed via the mean squared error (MSE), the spread-error relationship, and the CRPS. The predicted climate is assessed via a single long-time ensemble prediction, computed over 3000 MTU. The resulting distributions of the slow variables are compared via the Kolmogorov–Smirnov distance and the Hellinger distance. All error assessment metrics are further detailed in Appendix B. A summary of the models tested in this section is given in Table 3.
@@ 4.1 Short-time predictions
A further increase in average accuracy is observed for the CRPS-based models. Increasing the trajectory length generally leads to parametrizations with improved accuracy, up to a certain limit. For example, the additive noise model trained at $N_{t}=8$ is more accurate than that trained at $N_{t}=2$, and is approximately as accurate as that trained at $N_{t}=32$. The multiplicative models exhibit an even smaller MSE regardless of the trajectory length used in training, suggesting that the additional degrees of freedom allow for increased prediction accuracy.
@@
For the trained models, it could be desirable to diminish the weight of the self-spread term in the CRPS loss to possibly favor statistical consistency. This adaptation requires an additional hyperparameter or a term penalizing the discrepancy between the ensemble spread and MSE. Such a modification might diminish the ensemble sharpness during training without sacrificing accuracy. We investigate the use of a hyperparameter to scale the self-spread term in Section 4.4.
@@ 4.2 Long-time predictions
The climate prediction using the stochastic parametrizations is shown in Fig. 7, defined via the empirical PDFs of the slow variables over a long simulation time (3000 MTU). Here, the ‘crps_mult8’ is not included in the results, as it exhibited numerical instabilities during extended simulation. Generally, all stochastic parametrizations produce qualitatively similar symmetric PDFs with an excessively large standard deviation. An improvement of the CRPS-based models is found compared to the derivative-fitting-based models, which becomes clearer when increasing the trajectory length. This is summarised in Table 4, indicating that the additive noise models learned using longer trajectory lengths outperform the other tested models. However, the results are not as accurate as those presented in literature (see, e.g., [16], [10]), suggesting that combining CRPS-based learning with other state-dependent model forms might be beneficial for long-time prediction accuracy.
@@ 4.3 Short-time predictions with perturbed initial conditions
4.4 Sensitivity to self-spread penalty
The previous subsections highlight that the multiplicative noise models generate sharp ensemble forecasts, particularly visible at shorter lead times. This is a consequence of using the CRPS as the loss function, which balances accuracy and ensemble spread, favouring slightly dispersed forecasts that still reproduce the observed value during training. From a statistical perspective, however, forecast reliability requires that the ensemble spread equals the ensemble error [43, 31]. We therefore investigate the role of the spread term in the CRPS loss function and quantify its influence on short-term forecasts.
To this end, we modify the loss function (7) so that the self-spread penalty is scaled by a factor $\alpha$,
$\begin{split}\mathrm{Loss}(\theta,q,\alpha)&=\sum_{j=1}^{N_{t}}\mathrm{CRPS}(z_{\theta},C(q),t_{j},\alpha),\\ \mathrm{CRPS}(z_{\theta},C(q),t_{j},\alpha)&=\frac{1}{M}\sum_{i=1}^{M}\left|z_{\theta}^{i}(t_{j})-C(q(t_{j}))\right|-\frac{\alpha}{2M^{2}}\sum_{j=1}^{M}\sum_{k=1}^{M}\left|z_{\theta}^{i}(t_{j})-z_{\theta}^{k}(t_{j})\right|,\end{split}$ (19)
where $\alpha\in[0,1]$. We henceforth refer to (19) as the scaled CRPS loss function. The addition of $\alpha$ allows for training models that minimize the ensemble mean error while reducing the importance of the ensemble spread. The standard CRPS loss function (7) is recovered when $\alpha=1$. The forecast spread term is progressively diminished when $\alpha$ decreases, and vanishes entirely when $\alpha=0$, leaving only the ensemble mean error in the loss function. The value of $\alpha$ cannot exceed 1 to ensure nonnegativity of the scaled CRPS.
The effects of scaling the CRPS loss function are assessed by repeating the model training and forecasting procedures for a sequence of values of $\alpha$. A single model form is selected for the sake of clarity. Here, we adopt the multiplicative noise model and use a trajectory length $N_{t}=8$ to train the model, corresponding to the ‘crps_mult8’ setting considered in the previous subsections. This model was found to produce accurate yet sharp ensemble forecasts, and is therefore suited to demonstrate the ability to change the ensemble spread via choice of the loss function.
The evolution of the training loss is shown in Fig. 9, demonstrating consistent behaviour across the two test cases. Starting from the same initial loss value, the loss decreases at a similar rate for each values of $\alpha$. Smaller final losses are achieved for smaller $\alpha$. This is expected, since a reduced weight on the self-spread term in the scaled CRPS imposes fewer or less stringent constraints during training, thereby facilitating easier optimisation of the loss function.
Figure 9: Evolution of the loss per batch over the training duration, for $c=4$ (left panel) and $c=10$ (right panel), for several values of $\alpha$ in the scaled CRPS loss function (19). The displayed values are normalised by the initial value for visual comparison, and a moving average is applied for clarity.
A comparison between the ensemble error and ensemble spread is provided in Fig. 10. A clear effect on the ensemble spread is observed in both test cases, with the forecast sharpness increasing as $\alpha$ decreases. This occurs because, when $\alpha=1$, the CRPS formulation favours a larger spread to minimise the loss. At the same time, the ensemble error increases with decreasing $\alpha$. This suggests that retaining the self-spread term in the CRPS is beneficial for reducing mean error, despite not directly contributing to the error itself. Larger values of $\alpha$ regularise training by preventing a collapse of ensemble members, thereby reducing overfitting and generally yielding more stable predictions.
For the same reason, the spread observed for $c=10$ and small values of $\alpha$ initially remains small but later exceeds the spread obtained with $\alpha=1$. This behaviour is also attributed to the regularising property of the self-spread term in the loss function. Namely, models trained with small $\alpha$ are more sensitive to perturbations, and their forecast spread is therefore amplified more at later simulation times. These results demonstrate that the ensemble spread and model robustness can be systematically controlled through the choice of loss function, specifically via the weighting of the self-spread term.
Figure 10: Comparison of ensemble error (solid lines) and ensemble spread (dashed lines) for short lead times, for several values of $\alpha$ in the scaled CRPS loss function (19). The error and spread are computed per $X$-variable in ensembles consisting of 50 members, simulated from 100 different initial conditions.
@@ 5 Conclusion
In this paper, we investigated the calibration of stochastic parametrizations for ensemble forecasting using trajectory learning based on the continuous ranked probability score (CRPS). Using the two-scale Lorenz ’96 system as a testbed, we demonstrated that this approach yields models that provide both accurate and sharp short-term forecasts. Both additive and multiplicative noise models, leveraging global spatial patterns from POD modes, were found to outperform models based on derivative fitting. The improved accuracy at short lead times suggests that CRPS-based trajectory fitting is promising approach when combined with data assimilation, where accurate short-term forecasts are essential. Moreover, the self-spread term in the CRPS was found to have a stabilising effect on the model. Ultimately, this work establishes CRPS-based trajectory learning as a promising methodology for developing data-driven stochastic parametrizations for complex dynamical systems, representing a first step towards its application in geophysical flows.
@@ Appendix A Model and training details
The parameters for the derivative-fitting-based models with correlated noise, introduced in Section 3.2, are summarised in Table 5. These models use increments generated by an Ornstein–Uhlenbeck process, discretised as a first-order autoregressive (AR(1)) model, for the stochastic model component. The specific autocorrelation coefficients per POD mode are provided. Fig. 11 provides a visual comparison between the measured autocorrelation from the data and the autocorrelation function of the AR(1)-processes with fitted parameters. As discussed in Section 4.1, this figure highlights a potential overestimation of the temporal correlation.
@@ Appendix B Forecast assessment metrics
In the following definitions, we denote the truth by $y$ and denote by $x$ an ensemble forecast of size $M$, consisting of ensemble members $x_{i},\penalty 10000\ i=1,\ldots,M$. All $x_{j}$ and $y$ are regarded as independent identically distributed random variables with mean $\mu$ and variance $\sigma^{2}$.